Integrate x²dx/(9-x²)½ using trigonometric substitution
1
Expert's answer
2020-02-21T09:38:13-0500

"I=\int \frac{x^2}{\left(9-x^2\right)^{\frac{1}{2}}}dx"

Let "x = 3 sin(z)"

"\Rightarrow dx=3cos(z)dz"

substituting these value in "I"

"I=\int \frac{9sin^2\left(z\right)}{\left(9-9sin^2(z)\right)^{\frac{1}{2}}}\cdot 3cos\left(z\right)dz"

"=\int \frac{9sin^2\left(z\right)}{\left(9\left(1-sin^2\left(z\right)\right)\right)^{\frac{1}{2}}}\cdot 3cos\left(z\right)dz"

"=\int\frac{9sin^2\left(z\right)}{\left(9cos^2\left(z\right)\right)^{\frac{1}{2}}}\cdot 3cos\left(z\right)dz"

"=\int\frac{9sin^2\left(z\right)}{3cosz}\cdot 3cos\left(z\right)dz"

"=9\int sin^2(z)dz"

"=9 \int \:\frac{1-cos(2z)}{2}dz" as "cos(2z)=1-2sin^2(z)"

"=\frac{9}{2}\int \:\left(1-cos\left(2z\right)\right)dz"

"=\frac{9}{2}\left(z-\frac{sin\left(2z\right)}{2}\right)+C"

We assumed "x=3sin(z)"

"\Rightarrow" "sin(z)=\frac{x}3"

"z=sin^{-1}(\frac{x}{3})"

substituting back these value


"I=\frac{9}{2}\left(sin^{-1}\left(\frac{x}{3}\right)-\frac{1}{2}sin\left(2sin^{-1}\left(\frac{x}{3}\right)\right)\right)+C"

where C is the constant of integration.